MAT 142, HONORS CALCULUS, SUNY AT STONY BROOK, FALL 2005

Lecturer: Prof. Christopher Bishop Email: bishop@math.sunysb.edu Office: 4-112 in Math Building. Office hours: TBA.

Text: Calculus, 11th edition (silver cover), Thomas.

Webpages: Some wepages of interest: The math department wepage: www.math.sunysb.edu My wepage (has link to MAT 142 webpage): www.math.sunysb.edu/~bishop

TA’s webpage www.math.sunysb.edu/~craig

Syllabus: We will cover approximately Chapters 7,8,9 and 11 with certain omissions and additions of material. A more detailed syllabus will be provided.

Classdays. Classes begin on Monday, August 29 and end on Tuesday, December 13.

Holidays: There are no classes on Labor Day (Monday, September 5), Rosh Hashana (Mon-Wed, October 3-5), Yom Kipur (Thursday, October 13) and Thanksgiving (November 24-26).

Lectures times: Monday, Friday 12:50-2:10.

Lecture location: Room P-112 in the Physics building.

Recitation times and locations: There are two recitiations.

• Section 1 meets Tu-Th at 12:50-1:45 in Physics P-112. The TA is Craig vanCoevering. His office number is 2-104 in the Math Building and his email is craig@math.sunysb.edu. His office hours are TBA.

• Section 2 meets MW at 3:50-4:45 in ESS 177. Craig vanCoevering is also the TA for this section.

Homework. A list of problems from the text and supplementary material will be handed out. When a section is covered in lecture on Friday, the homework for that section is due at the first recitation of the following week. If it is covered on Monday, the homework is due by the second recitation of that week. The recitation instructor will collect and return homework. He may also give weekly quizzes.

Examinations. There will be two midterm tests: Friday, October 7 and Friday, November 11. These will be given during the regular lecture time. The final exam is on Friday, December 16 at 11:00am-1:30pm. The location will be announced in class during the last week of the semester.

Grades. Recitation 10%, Midterms 25% each, Final 40%.

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: This is in the basement of the Math Tower. It is staffed most days and some evenings; an exact schedule will be posted in the room. Note that this room is closed during final exam week. It also gets crowded before exams, so don’t wait until the last minute to get help there.

Calculus Online: There are many math sites on the internet. For information on the history of Calculus and short biographies of famous mathe-

maticians, try http://www-groups.dcs.st-and.ac.uk/~history and click

on History Topics or the Biography Index. You can find information about math classes at Stony Brook at www.math.sunysb.edu

Drop dates: Last day for Add/Drop is Sunday, September 4. Last day to switch to lower numbered math course is Friday, October 21. The last day to withdraw from a course is Friday, October 28.